OpenStudy (anonymous):

At the GAS n GO station, a car wash costs $4 with any gas purchase. Gas is selling at $2.38 per gallon. George can spend at least $30 but no more than $40 for gas and a car wash. How many gallons of gas can he buy?

OpenStudy (anonymous):

\[40\ge2.38x+4\ge 30\]\[40-4\ge2.38x \ge30-4\]\[\frac{36}{2.38}\ge x \ge \frac{26}{2.38}\]\[15\ge x \ge 11\]click good answer , if you like my answer,ty